OpenStudy (timaashorty):

what is the graph of the equation ? (x+8)^2+(y-1)^2=9? Please show your work.

OpenStudy (anonymous):

Looks like a circle.

OpenStudy (timaashorty):

\[(x+8)^{2} + (y-1)^{2} = 9\]

OpenStudy (timaashorty):

I know it's a circle and how it looks graphed. I just need to know before I graph it, do I need to do any work.

OpenStudy (shamil98):

(x-h)^2 + (y-k)^2 = r^2 r = 3 radius of circle (h,k) = the center of the circle.

OpenStudy (timaashorty):

(8, 1) is the center?

OpenStudy (shamil98):

no, (-8,1) is.

OpenStudy (timaashorty):

Oops I see. can you show me how you got the radius

OpenStudy (shamil98):

(x-h)^2 + (y-k)^2 = r^2 (x-(-8)^2 + (y-(+1)^2 = 3^2

OpenStudy (shamil98):

you have 9, r^2 = 9 r = 3

OpenStudy (timaashorty):

Oh Okay  just to make sure, This is how it will be graphed right?

OpenStudy (anonymous):

That's a little sloppy, but it gets the point across. If you label some points with their coordinates it would be more convincing.

OpenStudy (shamil98):

Yes, but the radius is 3 so remember the distance between the circle's points are 3.

OpenStudy (timaashorty):

yeah i know, it's just im in a rush right now, but it's similar?

OpenStudy (timaashorty):

Thanks for the help guys